OpenStudy (anonymous):

simplify 4-8*3/4+1/4 A. -7/4 B. -3 C. 11/4 D. 3

OpenStudy (anonymous):

You need to put some brackets for me to understand that. That can be expressed in many forms. Use brackets so i can determine the proper way to simplify that.

OpenStudy (anonymous):

sorry 4-8*(3/4)+(1/4)

OpenStudy (anonymous):

Okay. \[4-8(3/4) + (1/4)\] Expand the middle term. \[=4 -(24/4) + (1/4)\] \[= 4-6 + (1/4)\] \[= -2+ (1/4)\] \[= - (7/4)\] Therefore A is the answer.